OpenStudy (anonymous):

Carlos buys a box of frozen flavored ice sticks. \frac{2}{5}of the ice sticks are blue raspberry. Of the remaining ice sticks, \frac{1}{3}are grape, and the rest are cherry. If there are 48 blue raspberry and cherry flavored ice sticks altogether, how many ice sticks are in the box? 60 ice sticks 90 ice sticks 120 ice sticks 340 ice sticks

OpenStudy (anonymous):

can u hurry?

OpenStudy (texaschic101):

2/5 are blue raspberry of the remaining ice sticks (which there are 3/5 remaining)...1/3 are grape. 1/3 of 3/5 = 1/3 * 3/5 = 3/15....so 3/15 are grape. and 3/5 - 3/15 = 9/15 - 3/15 = 6/15...these are cherry so 12/15 are blue raspberry and cherry..... 48 is 12/15 of what ? 48 = 12/15x 48/(12/15) = x 48 * 15/12 = x 60 = x check.. 2/5 are blue raspberry... 2/5 of 60 = 2/5 * 60 = 120/5 = 24 are blue raspberry 3/15 are grape.. 3/15 of 60 = 3/15 * 60 = 180/15 = 12 are grape 6/15 are cherry... 6/15 of 60 = 6/15 * 60 = 360/15 = 24 are cherry 24 + 24 + 12 = 60 48 + 12 = 60 60 = 60 (correct) there are 60 ice sticks in the box

OpenStudy (anonymous):

thanks could u help me with 1 more?

OpenStudy (anonymous):

Of the books Joe sorted in the library, \frac{1}{3}were fiction. After he sorted out the fiction books, there were 136 books left. How many books did Joe have to sort at the start? 105 books 119 books 144 books 204 books

OpenStudy (texaschic101):

if 1/3 were fiction, then 2/3 were something else 2/3 of what = 136 2/3x = 136 x = 136/(2/3) x = 136 * 3/2 x = 408/2 x = 204 check... 1/3 were fiction.. 1/3 of 204 = 1/3 * 204 = 204/3 = 68 were fiction 2/3 were not fiction.. 2/3 of 204 = 2/3 * 204 = 408/3 = 136 were not fiction 136 + 68 = 204 204 = 204 (correct) Joe had 204 books to sort

OpenStudy (anonymous):

u r u good freind

OpenStudy (texaschic101):

thanks :)

OpenStudy (anonymous):

i sent u a message